A flying pig with a mass of 0.05 kg has elastic arms that have a combined k value of 750 N/m. If the arms are stretched by 25 cm, what speed will the pig have at the point of 0 cm stretch?


What height will the pig reach if it is launched straight up?


1
Expert's answer
2021-02-17T10:41:49-0500

According to energy conservation:


"\frac12 kx^2=\frac12 mv^2,\\\space\\
v=x\sqrt{\frac km}=30.6\text{ m/s}."

The height:


"h=\frac{ v^2}{2g}=47.8\text{ m}."